### v1

@@
22institutetext: Hamburger Sternwarte, Gojenbergsweg 112, 21029, Hamburg, Germany 33institutetext: University Observatory Munich, Faculty of Physics, Ludwig-Maximilians-Universität, Scheinerstr. 1, 81679, Munich, Germany 44institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstrasse 25/8, 6020, Innsbruck, Austria 55institutetext: Arnold Sommerfeld Center for Theoretical Physics, LMU Munich, Theresienstr. 37, 80333 Munich, Germany 66institutetext: INAF, Osservatorio Astronomico di Brera, Via E. Bianchi 46, 23807, Merate, (LC), Italy 77institutetext: Argelander Institute for Astronomy, University of Bonn, Auf dem Hügel 71, 53121, Bonn, Germany
The SRG/eROSITA all-sky survey.
@@
X. Zhang The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    E. Bulbul The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    N. Malavasi The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    V. Ghirardini The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    J. Comparat The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    M. Kluge The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    A. Liu The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    A. Merloni The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    Y. Zhang The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    Y. E. Bahar The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    E. Artis The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    J. Sanders The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    C. Garrel The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    F. Balzer The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    M. Brüggen The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    M. Freyberg The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    E. Gatuzz The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    S. Grandis The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    S. Krippendorf The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    K. Nandra The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    G. Ponti The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    M. Ramos-Ceja The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    P. Predehl The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    T. H. Reiprich The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    A. Veronica The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    M. C. H. Yeung The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.    S. Zelmer The SRG/eROSITA all-sky survey.The SRG/eROSITA all-sky survey.
@@
The properties of the warm-hot intergalactic medium (WHIM) in the cosmic filaments are among the least quantified in modern astrophysics. eROSITA All Sky Survey (eRASS) provides us with a unique opportunity to study the X-ray emission of the WHIM. We applied both imaging and spectroscopic stacking techniques to the data of the first four eRASS scans to inspect the X-ray emissions from 7817 cosmic filaments identified from Sloan Digital Sky Survey optical galaxy samples. We obtained a $9\sigma$ significant detection of the total X-ray signal from filaments in the 0.3–1.2 keV band. We introduce a novel method to estimate the contamination fraction from unmasked X-ray halos, active galactic nuclei, and X-ray binaries that are associated with galaxies in filaments. We found an approximately 40% contamination fraction by these unmasked sources, suggesting that the remaining 60% of the signal could be from the WHIM and a $5.4\sigma$ detection significance of the WHIM. Moreover, we modeled the temperature and baryon density contrast of the detected WHIM by fitting the stacked spectrum and surface brightness profile. The best-fit temperature $\log(T/\mathrm{K})=6.84\pm 0.07$, obtained by using a single temperature model is marginally higher than the simulation results, which could be due to the fitting of a single temperature model on a multi-temperature spectrum. The best-fit baryon density contrast $\log\Delta_{\mathrm{b}}=1.88\pm 0.18$ generally agrees with the X-ray emitting phases in the IllustrisTNG simulation, suggesting that the broadband X-ray emission traces the high temperature and density end of the entire WHIM population.
@@ 1 Introduction
The warm-hot intergalactic medium (WHIM) permeating the cosmic filaments with temperature of $10^{5-7}$ K and density $n_{\mathrm{H}}<10^{-4}$ cm-3 is the majority of the “missing baryon” in the low redshift Universe predicted by numerical simulations (e.g., Fukugita et al. 1998; Cen & Ostriker 1999; Shull et al. 2012; Tuominen et al. 2021). Recent measurements on the dispersions of localized fast radio bursts that trace the total line-of-sight free electron column density quantify the cosmic baryon density in the low redshift Universe in an new way (Macquart et al. 2020; Yang et al. 2022; Wang & Wei 2023). These FRB cosmic baryon density measurements are consistent with the results from other probes such as cosmic microwave background (CMB) (Planck Collaboration et al. 2016, 2020) and Big Bang Nucleosynthesis (BBN) (Pitrou et al. 2018), which conclude that baryons are not missing in the low redshift Universe. However, the FRB dispersion measure technique is unable to provide additional detailed information about the baryons such as density and temperature. Exploring these properties by using other observation techniques such as X-ray observations and Sunyaev–Zeldovich (SZ) effects is still one of the ongoing prime investigations of modern astrophysics (Driver 2021).
@@
Currently, absorption studies by using OVI lines in the far ultraviolet bands can successfully trace the $10^{5.5}$ K low-temperature phases WHIM (Danforth et al. 2016). For X-ray absorption lines that trace $\sim 10^{6}$ K phases WHIM, though there are many cases of plausible detections (Nicastro et al. 2005; Fang et al. 2010; Nicastro et al. 2018), these constraints suffer from potential contamination from other absorption sources (e.g. Johnson et al. 2019; Gatuzz et al. 2023). This method is also currently restricted to a handful of bright background blazars (see Fang et al. 2022 for a review). Meanwhile, X-ray emission studies have been mostly focused on the brightest part of the WHIM with relatively high-density contrast, which resides in bridges connecting galaxy cluster pairs or along the filaments (Werner et al. 2008; Bulbul et al. 2016; Akamatsu et al. 2017; Reiprich et al. 2021) or circumcluster space of massive galaxy clusters (Eckert et al. 2015; Bulbul et al. 2016; Reiprich et al. 2021). Similarly, the number of these systems is limited, and there is less information about the properties of the WHIM in the environments of long cosmic filaments.
@@
Wide-area optical spectroscopic surveys have mapped the large-scale galaxy distributions up to redshifts of 0.7 (Reid et al. 2016), offering opportunities for a complete examination of the WHIM emission in the vast space of the cosmic filaments. Two pilot studies have shown mild positive detections ($4.2\sigma$ and $3.8\sigma$) of stacked X-ray surface brightness emission from cosmic filaments using ROSAT All-Sky Survey (RASS) and Final Equatorial-Depth Survey (eFEDS) data (Tanimura et al. 2020b, 2022), demonstrating the potential of X-ray stacking analysis to trace the emission in cosmic filaments. Similarly, mild positive detections have also been reported by stacking the SZ effect signal that traces the pressure of the intergalactic medium (de Graaff et al. 2019; Tanimura et al. 2019, 2020a). However, in the previous work, the detection in the very soft X-ray band $<0.6$ keV at the observed frame, which is needed to constrain the temperature of the redshifted $T<10^{7}$ K gas, is still lacking. Meanwhile, The energy resolution was limited in the observer frame due to the broad redshift distribution of the filament sample, limiting the ability to measure the plasma temperature in the soft band. Moreover, the pilot X-ray image stacking studies did not account for the signal from unmasked X-ray sources, which could contribute a significant fraction of the total emission. In this work, we improve upon the previous stacking analyses by extending the bandpass for spectral analysis, increasing the energy resolution by using a rest-frame spectral stacking scheme, accounting for the emission from unmasked low-mass halos and point sources, and utilizing the new All-Sky Survey (eRASS) data that covers an extensive area with an unprecedented depths, with excellent soft band sensitivity that is essential for detecting the faint X-ray emission from WHIM gas.
@@
In this work, we assume a fiducial Lambda cold dark matter cosmology with parameters $H_{0}=70$ km s-1 Mpc-1, $\Omega_{\mathrm{m}}=0.3$, $\Omega_{\Lambda}=0.7$, and adopt the ratio between baryon density parameter to matter density parameter from Planck18 cosmology results, which is $\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}=0.158$ (Planck Collaboration et al. 2020). At $z=0$, the mean baryon density is $\overline{\rho}_{\mathrm{b}}(0)=4.51\times 10^{-31}$ g cm-3. The baryon density contrast is defined as $\Delta_{\mathrm{b}}\equiv\rho_{\mathrm{b}}/\overline{\rho}_{\mathrm{b}}$. Throughout this article, $\log$ refers to the logarithm to the base 10, and $\ln$ refers to the natural logarithm. Without further indication, all sizes/scales/lengths are in physical/proper scale. This article is organized as follows. In Sect. 2, we describe the X-ray data products and the optical filament catalog we use. In Sect. 3 we present the surface brightness profile stacking method and results. In Sect. 4, we estimate the fraction of detected signal due to unmasked galactic sources. Sect. 5 presents the spectral stacking method and temperature measurements. The estimation of the detected X-ray emitting phases and the comparison with numerical simulations will be presented in Sect. 6. We conclude our work in Sect. 7.
@@ 2 Observations, optical filament catalog, and data preparation
The stacked first four scans of eROSITA All-Sky Survey (hereafter eRASS:4) data were collected from 12th December 2019 to 19th December 2021. The data was processed with eSASS (Brunner et al. 2022) pipeline version 020, which is the same as the version 010 used for eRASS1 data release with improvements on boresight correction, detector noise suppression, and subpixel position computation (Merloni et al. 2024). We only select the events from telescope modules 1,2,3,4,6 (hereafter TM8) to avoid the systematic uncertainties caused by the optical light leak in TMs 5 and 7 (Predehl et al. 2021). We used the tools in the software package eSASS version eSASSusers_211214_0_4 to generate eRASS data products.
We adopt an optical filament catalog (Malavasi et al. 2020) compiled using the DisPerSE software (Sousbie 2011; Sousbie et al. 2011) based on the Sloan Digital Sky Survey (SDSS) LOWZ+CMASS galaxy catalogs (Reid et al. 2016). The filament catalog contains 63391 filaments after a $3\sigma$ persistence cut without any smoothing applied on the galaxy density field. The original filament catalog is stored in the comoving frame and Cartesian coordinates. Therefore, we convert the coordinates back to (Right Ascension, Declination, redshift) using the conversion functions and the cosmology used in Malavasi et al. (2020). The physical length of filaments is calculated using the assumed cosmological parameters above. The cross-section of the SDSS LOWZ+CMASS footprint (Reid et al. 2016) and the Western Galactic hemisphere of the eROSITA All-Sky Survey determines a $\sim 3200\deg^{2}$ common analysis footprint. In addition, the bright foreground regions of the eROSITA Bubble (Predehl et al. 2020), the Virgo Cluster (Böhringer et al. 1994; McCall et al. 2024), and the Galactic plane $|b|<20^{\circ}$ are excluded from the analysis for simplicity. The total analysis footprint before source masking has a solid angle shown in Fig. 1, amounting to 2275 deg2. In this footprint, we select 7817 filaments in a redshift range $0.60.2<z<$ 0.6 and a physical length range 20 Mpc $<l_{\mathrm{fil}}<$ 100 Mpc for stacking analysis (see Fig. 2).

### v2

@@
22institutetext: INAF, Osservatorio di Astrofisica e Scienza dello Spazio, via Piero Gobetti 93/3, 40129 Bologna, Italy 33institutetext: Hamburger Sternwarte, Gojenbergsweg 112, 21029, Hamburg, Germany 44institutetext: University Observatory Munich, Faculty of Physics, Ludwig-Maximilians-Universität, Scheinerstr. 1, 81679, Munich, Germany 55institutetext: Universität Innsbruck, Institut für Astro- und Teilchenphysik, Technikerstrasse 25/8, 6020, Innsbruck, Austria 66institutetext: Arnold Sommerfeld Center for Theoretical Physics, LMU Munich, Theresienstr. 37, 80333 Munich, Germany 77institutetext: INAF, Osservatorio Astronomico di Brera, Via E. Bianchi 46, 23807, Merate, (LC), Italy 88institutetext: Argelander Institute for Astronomy, University of Bonn, Auf dem Hügel 71, 53121, Bonn, Germany
The SRG/eROSITA all-sky survey
@@
X. Zhang The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    E. Bulbul The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    N. Malavasi The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    V. Ghirardini The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    J. Comparat The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    M. Kluge The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    A. Liu The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    A. Merloni The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    Y. Zhang The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    Y. E. Bahar The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    E. Artis The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    J. S. Sanders The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    C. Garrel The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    F. Balzer The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    M. Brüggen The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    M. Freyberg The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    E. Gatuzz The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    S. Grandis The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    S. Krippendorf The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    K. Nandra The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    G. Ponti The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    M. Ramos-Ceja The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    P. Predehl The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    T. H. Reiprich The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    A. Veronica The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    M. C. H. Yeung The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey    S. Zelmer The SRG/eROSITA all-sky surveyThe SRG/eROSITA all-sky survey
@@
The properties of the warm-hot intergalactic medium (WHIM) in cosmic filaments are among the least quantified units in modern astrophysics. The Spectrum Roentgen Gamma/eROSITA All Sky Survey ((SRG/eRASS) provides a unique opportunity to study the X-ray emission of the WHIM. We applied both imaging and spectroscopic stacking techniques to the data of the first four eRASS scans to inspect the X-ray emissions from 7817 cosmic filaments identified from Sloan Digital Sky Survey (SDSS) optical galaxy samples. We obtained a $9\sigma$ significant detection of the total X-ray signal from filaments in the 0.3–1.2 keV band. Here, we introduce a novel method to estimate the contamination fraction from unmasked X-ray halos, active galactic nuclei, and X-ray binaries associated with filament galaxies. We found an approximately 40% contamination fraction for these unmasked sources, suggesting that the remaining 60% of the signal could be coming from the WHIM and a $5.4\sigma$ detection significance of the WHIM. Moreover, we modeled the temperature and baryon density contrast of the detected WHIM by fitting the stacked spectrum and surface brightness profile. The best-fit temperature $\log(T/\mathrm{K})=6.84\pm 0.07$, obtained by using a single temperature model, is marginally higher than in the simulation results. This could be due to the fitting of a single temperature model on a multi-temperature spectrum. Assuming a 0.2 solar abundance, the best-fit baryon density contrast $\log\Delta_{\mathrm{b}}=1.88\pm 0.18$ is in general agreement with the X-ray emitting phases in the IllustrisTNG simulation. This result suggests that the broadband X-ray emission traces the high end of the temperature and density values that characterize the entire WHIM population.
@@ 1 Introduction
The warm-hot intergalactic medium (WHIM) permeating cosmic filaments, with a temperature of $10^{5-7}$ K and density of $n_{\mathrm{H}}<10^{-4}$ cm-3, makes up the majority of the “missing baryon” in the low-redshift Universe predicted by numerical simulations (e.g., Fukugita et al. 1998; Cen & Ostriker 1999; Shull et al. 2012; Tuominen et al. 2021). Recent measurements on the dispersions of localized fast radio bursts that trace the total line-of-sight free electron column density quantify the cosmic baryon density in the low-redshift Universe in a new way (Macquart et al. 2020; Yang et al. 2022; Wang & Wei 2023). These FRB cosmic baryon density measurements are consistent with the results from other probes, such as cosmic microwave background and big bang nucleosynthesis (Pitrou et al. 2018), which conclude that baryons are not missing in the low-redshift Universe. However, the FRB dispersion measure technique cannot provide additional detailed information on the baryons, such as density and temperature. Exploring these properties using other observation techniques, such as X-ray observations and Sunyaev–Zeldovich (SZ) effects, remains one of the primary investigation goals in modern astrophysics (Driver 2021).
@@
Currently, absorption studies by using OVI lines in the far ultraviolet bands can successfully trace the $10^{5.5}$ K low-temperature phases WHIM (Danforth et al. 2016). For X-ray absorption lines that trace $\sim 10^{6}$ K phases WHIM, while there are many cases of plausible detections (Nicastro et al. 2005; Fang et al. 2010; Nicastro et al. 2018), these constraints suffer from potential contamination from other absorption sources (e.g. Johnson et al. 2019; Dorigo Jones et al. 2022; Gatuzz et al. 2023). This method is also currently restricted to a handful of bright background blazars (see Fang et al. 2022 for a review). Meanwhile, X-ray emission studies have been mostly focused on the brightest part of the WHIM with a relatively high density contrast, which resides in bridges connecting galaxy cluster pairs or along the filaments (Werner et al. 2008; Bulbul et al. 2016; Akamatsu et al. 2017; Reiprich et al. 2021) or circumcluster space of massive galaxy clusters (Eckert et al. 2015; Bulbul et al. 2016; Reiprich et al. 2021). Similarly, the number of these systems is limited, and there is less information about the properties of the WHIM in the environments of long cosmic filaments.
@@
Wide-area optical spectroscopic surveys have mapped the large-scale galaxy distributions up to redshifts of 0.7 (Reid et al. 2016), thus offering opportunities for a complete examination of the WHIM emission in the vast space of the cosmic filaments. Two pilot studies have shown mild positive detections ($4.2\sigma$ and $3.8\sigma$) of stacked X-ray surface brightness emission from cosmic filaments using ROSAT All-Sky Survey (RASS) and Final Equatorial-Depth Survey (eFEDS) data (Tanimura et al. 2020b, 2022), demonstrating the potential of X-ray stacking analysis to trace the emission in cosmic filaments. Similarly, mild positive detections have also been reported by stacking the SZ effect signal that traces the pressure of the intergalactic medium (de Graaff et al. 2019; Tanimura et al. 2019, 2020a). However, in these previous works, detections in the very soft X-ray band $<0.6$ keV at the observed frame, needed to constrain the temperature of the redshifted $T<10^{7}$ K gas, are still lacking. Meanwhile, the energy resolution was limited in the observer frame due to the broad redshift distribution of the filament sample, limiting the ability to measure the plasma temperature in the soft band. Moreover, the pilot X-ray image stacking studies did not account for the signal from unmasked X-ray sources, which could contribute a significant fraction of the total emission. In this work, we improve upon the previous stacking analyses by extending the bandpass for spectral analysis, increasing the energy resolution by using a rest-frame spectral stacking scheme, accounting for the emission from unmasked low-mass halos and point sources and utilizing the new All-Sky Survey (eRASS) data. This gives us access to an extensive area with an unprecedented depths, with excellent soft band sensitivity that is essential for detecting the faint X-ray emission from WHIM gas.
@@
In this work, we assume a fiducial Lambda cold dark matter cosmology with parameters: $H_{0}=70$ km s-1 Mpc-1, $\Omega_{\mathrm{m}}=0.3$, $\Omega_{\Lambda}=0.7$. We adopt the ratio between baryon density parameter to matter density parameter from the Planck18 cosmology results, namely: $\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}=0.158$ (Planck Collaboration et al. 2020). At $z=0$, the mean baryon density is $\overline{\rho}_{\mathrm{b}}(0)=4.51\times 10^{-31}$ g cm-3. The baryon density contrast is defined as $\Delta_{\mathrm{b}}\equiv\rho_{\mathrm{b}}/\overline{\rho}_{\mathrm{b}}$. Throughout this article, $\log$ refers to the logarithm to the base 10, and $\ln$ refers to the natural logarithm. Unless otherwise indicated, all sizes, scales, and lengths are given in physical (proper) scale. This article is organized as follows. In Sect. 2, we describe the X-ray data products and the optical filament catalog we used. In Sect. 3, we present the surface brightness profile stacking method and results. In Sect. 4, we estimate the fraction of detected signal due to unmasked galactic sources. Section 5 presents the spectral stacking method and temperature measurements. The estimation of the detected X-ray emitting phases and the comparison with numerical simulations are presented in Sect. 6. We conclude our work in Sect. 7.
@@ 2 Observations, optical filament catalog, and data preparation
The stacked first four scans of the eROSITA All-Sky Survey (hereafter eRASS:4) data were collected from December 12, 2019 to December 19, 2021. The data were processed with eSASS (Brunner et al. 2022) pipeline version 020, which is the same as the version 010 used for eRASS1 data release with improvements on boresight correction, detector noise suppression, and subpixel position computation (Merloni et al. 2024). We only selected events from telescope modules 1, 2, 3, 4, and 6 (hereafter, TM8) to avoid the systematic uncertainties caused by the optical light leak in TMs 5 and 7 (Predehl et al. 2021). We used the tools in the software package eSASS version eSASSusers_211214_0_4 to generate eRASS data products.
We adopted an optical filament catalog (Malavasi et al. 2020), compiled using the DisPerSE software (Sousbie 2011; Sousbie et al. 2011) and based on the Sloan Digital Sky Survey (SDSS) LOWZ+CMASS galaxy catalogs (Reid et al. 2016). The filament catalog contains 63391 filaments after a $3\sigma$ persistence cut without any smoothing applied on the galaxy density field. The original filament catalog is stored in the comoving frame and Cartesian coordinates. Therefore, we converted the coordinates back to (right ascension, declination, redshift) using the conversion functions and the cosmology used in Malavasi et al. (2020). The physical length of the filaments was calculated using the assumed cosmological parameters above. The cross-section of the SDSS LOWZ+CMASS footprint (Reid et al. 2016) and the western Galactic hemisphere of the eROSITA All-Sky Survey determines a $\sim 3200\deg^{2}$ common analysis footprint. In addition, the bright foreground regions of the eROSITA Bubble (Predehl et al. 2020), the Virgo Cluster (Böhringer et al. 1994; McCall et al. 2024), and the Galactic plane $|b|<20^{\circ}$ were excluded from the analysis for simplicity. The total analysis footprint before source masking has a solid angle is shown in Fig. 1, amounting to 2275 deg2. In this footprint, we select 7817 filaments in a redshift range $0.60.2<z<$ 0.6 and a physical length range 20 Mpc $<l_{\mathrm{fil}}<$ 100 Mpc for the stacking analysis (see Fig. 2).